OpenStudy (anonymous):

write the linear equation y+3=(2/5)(x+9) in standard form

OpenStudy (saifoo.khan):

Damn ! missed the x.

OpenStudy (saifoo.khan):

\[y + 3 = \frac25x + \frac25(9)\] \[y + 3 = \frac25x + \frac{18}{5}\] \[y + 3= \frac{2x+18}{5}\]

OpenStudy (saifoo.khan):

5(y+3) = 2x + 18

OpenStudy (saifoo.khan):

\[5y + 15 = 2x + 18\]

OpenStudy (saifoo.khan):

Im sure you can solve it now..

OpenStudy (anonymous):

na

OpenStudy (saifoo.khan):

Damn! c'mon man!

OpenStudy (saifoo.khan):

\[5y = 2x + 18 - 15\] \[5y = 2x + 3\]

OpenStudy (saifoo.khan):

\[y = \frac25x + \frac35\]

OpenStudy (saifoo.khan):

here you go!

OpenStudy (anonymous):

2x-5y=-3